a block a mass 50kg slides down a plane inclined at an angle 30 to the horizontal. if the coefficient of friction is 0.4 calculate the frictional force btwn the two surfaces (g=10m/s)
1
Expert's answer
2019-10-30T12:30:34-0400

The friction force by definition


"F=\mu N"

The normal reaction force is given by

"N=mg\cos\theta"

So


"F=\mu mg\cos\theta"

"=0.4\times 50\:\rm kg\times 10\: m/s^2\times \cos 30^{\circ}=173.2\:\rm N"